# Can In-context Learning Really Generalize to Out-of-distribution Tasks?

### v1

@@ 1 Introduction
Recent studies have demonstrated that ICL can tackle seemingly new tasks. For instance, Garg et al. (2022); Raventós et al. (2024); Zhang et al. (2023); Akyürek et al. (2022) found that ICL can learn new linear regression weights after pretraining on a large set of weight vectors. Moreover, Pan (2023); Kossen et al. (2024); Vacareanu et al. (2024) revealed that real-world LLMs like Llama-2 (Touvron et al., 2023) and GPT-4 (Achiam et al., 2023) are capable of solving artificially constructed tasks likely unseen in their pretraining data, such as a classification task with abstract labels.
@@
In this work, we aim to consolidate previous findings by addressing these questions. First, we empirically show that when pretrained on a specific function class, the OOD performance of ICL approaches that of a model from the same function class optimized via gradient descent. This suggests that ICL tends to implement functions encountered during pretraining, which could explain its failure on OOD tasks that significantly deviate from the training distribution. Furthermore, we reproduce the widely observed phenomenon that ICL can perform classification with abstract labels. We find that solving such tasks requires retrieving similar labels from the context, a capability that can be acquired through pretraining on analogous tasks. This implies that success in such tasks of ICL may not indicate an inherent ability to learn new tasks. Finally, we explore scenarios in which the model is pretrained on multiple tasks, empirically uncovering the algorithm selection mechanism for OOD tasks. Building on the work of Lin & Lee (2024), we also provide a comprehensive theoretical framework for understanding the ICL mechanism.
Our contributions are summarized as follows:
@@
    We empirically show that ICL tends to implement the pretraining function based on the downstream task context, highlighting its limitation in solving OOD tasks (Section 2).
@@
    We further investigate ICL’s ability to perform classification with unseen abstract labels. Although this appears to be evidence of ICL learning OOD tasks, we find that such tasks can be solved by retrieving similar examples from the context. This retrieval ability can arise from training on tasks with more diverse abstract labels (Section 3.1) and only emerges when the testing function is in distribution (Section 3.2). Additionally, we find that a pretrained Llama2-7b fails to learn OOD functions through ICL in a synthetic word classification task (Section 3.3), further confirming ICL’s limitations in OOD scenarios.
@@
2 Transformers Implements Functions Classes Seen During Pretraining
@@
Exploring the ICL performance on OOD tasks. To investigate the ICL performance on new tasks that are unseen during training, following Garg et al. (2022), we train a GPT-2 (Radford et al., 2019) from scratch on some simple functions and evaluate it on functions different from the training ones. Denote the Transformer model parameterized by $\theta$ as $M_{\theta}$. The pretraining objective is:
$\min_{\theta}\frac{1}{T}\sum_{i=1}^{T}\mathbb{E}_{f\sim\mathcal{F}}[\|M_{\theta}(\mathcal{S}_{i}\oplus\bm{x}_{i+1})-f(\bm{x}_{i})\|^{2}_{2}],$ (1)
@@
where $\mathcal{S}_{i}=[\bm{x}_{1}\oplus y_{1}\oplus\bm{x}_{2}\oplus y_{2}\oplus...\oplus\bm{x}_{i}\oplus y_{i}]\in\mathbb{R}^{d\times 2i}$ is the context of length $i$, $\oplus$ denotes concatenation. $\bm{x}_{i}\in\mathbb{R}^{d}$ are sampled from a standard normal distribution $\mathcal{N}(0,1)$ with dimension $d=20$. Let $y_{i}=f(\bm{x}_{i})$ represent the labels, with $\mathcal{F}$ denoting the hypothesis class to which $f$ belongs. We train three separate GPT-2 models on three different function classes $\mathcal{F}$: linear regression, quadratic regression (element-wise square followed by linear regression), and a 2-layer ReLU network (detailed descriptions are in Appendix A.1). We then evaluate their ICL performance on these three tasks. Note that even when the testing and training functions are i.i.d. sampled from the same task, the specific instances of the testing functions remain unseen during training. For comparison, we also train models within the corresponding $\mathcal{F}$ with gradient descent (GD) using the testing in-context examples (details in Appendix A.1).
@@
Observations. We plot the test error on the three tasks in Figure 1 and observe that: 1) when evaluated on the same task $\mathcal{F}$ as pretraining, ICL can reach near-zero test error, as observed in previous works (Garg et al., 2022). 2) when evaluated on a new task, ICL performs similarly to the corresponding model of the pretraining function class optimized by GD, given enough in-context examples. This indicates that the ICL prediction implicitly implements function classes seen during pretraining. 3) The models trained on linear and quadratic regression exhibit a double descent error curve (Nakkiran, 2019), characterized by a high error when given exact $d$ examples and evaluated on a new task, which has been theoretically and empirically revealed under the noisy linear regression setting by Nakkiran (2019) and Garg et al. (2022), respectively. This further demonstrates that ICL implements the linear regression pretraining task, as the double descent curve is a distinctive phenomenon unique to linear regression models.
@@
Figure 1: The ICL test error of Transformers trained on different function classes (solid lines) and the performance of the models from the corresponding pretraining functions classes trained by gradient descent (GD) using the in-context examples (dashed lines). Y-axis: test square error. X-axis: context length. The evaluation tasks are (a) linear regression, (b) quadratic regression, and (c) 2-layer ReLU network regression. In all sub-figures, we observe that as the test context length increases, the ICL performance of the Transformer pretrained on a particular function class closely approaches that of the model from this function class trained by GD.
Existing theoretical evidence. One recent work (Zhang et al., 2023) theoretically proved that a one-layer linear self-attention model (LSA, defined in Appendix B.1) pretrained on a linear regression task will still implement the linear predictor given downstream in-context examples of arbitrary new function classes, under some assumptions on the initialization of the Transformer weight matrices. We restate the Theorem 4.2 of Zhang et al. (2023) as Lemma 2.1 below:
Lemma 2.1.
(Theorem 4.2 of Zhang et al. (2023), informal) Let $\mathcal{D}$ be a distribution over $(\bm{x},y)\in\mathbb{R}^{d}\times\mathbb{R}$, whose marginal distribution on $x$ is $\mathcal{D}_{x}=$ $\mathcal{N}(0,\Lambda)$. Assume the test prompt is of the form $P=\left(\bm{x}_{1},y_{1},\ldots,\bm{x}_{T},y_{T},\bm{x}_{\text{query }}\right)$, where $\left(\bm{x}_{i},y_{i}\right),\left(\bm{x}_{\text{query }},y_{\text{query }}\right)\stackrel{{\scriptstyle\text{ i.i.d. }}}{{\sim}}\mathcal{D}$. The prediction risk on the test query $y_{query}$ of an arbitrary task satisfies:
Lemma 2.1 serves as a shred of theoretical evidence that ICL can just implement the pretraining function class, while the role of the context examples is to optimize the model within the pretraining hypothesis space. We leave Assumption B.1 that Lemma 2.1 depends on in Appendix B.1.
3 Learning Abstract Labels May Not Be A New-task-learning Ability
@@ 3.1 Classification Tasks with Unseen Abstract Labels
Recent works (Pan et al., 2023; Kossen et al., 2024) have shown that LLMs can perform classification tasks in which the labels are ”abstract symbols” with no semantic meaning. For instance, in the SST-2 binary classification task, the labels ’positive’ and ’negative’ are substituted with abstract terms like ’foo’ and ’bar’, respectively. These tasks are likely not seen during pretraining. Pan et al. (2023) refer to this ability of ICL to perform such classification as ”task learning” (TL). In this section, we explore whether the TL ability really reflects a new-task-learning capability of ICL or if it merely stems from the model having learned similar tasks during pretraining.
The retrieval ability can be gained by pretraining on a retrieval task with diverse input-label mappings. The classification of abstract labels can be approached by first retrieving an example with semantics similar to the query and then outputting the label of that example, as empirically demonstrated in previous research (Wang et al., 2023; Yu & Ananiadou, 2024). Therefore, the retrieval ability is a crucial prerequisite for performing abstract-label classification. To investigate whether ICL’s retrieval capability can emerge from training on similar tasks, we construct a retrieval task. Specifically, we generate a predefined word embedding $E\in\mathbb{R}^{N\times d}$ and randomly sample $\bm{x}_{i}\in\mathbb{R}^{d}$ from the first 5 rows of $E$. Let the index of $\bm{x}_{i}$ in $E$ be $I_{\bm{x}_{i}}$ (where $I_{x_{i}}\in[0,5)$). The labels $y_{i}$ are generated by mapping $I_{\bm{x}_{i}}$ to new indices $I_{y_{i}}\in[N]$ according to a mapping rule $I_{y_{i}}=R_{s}(I_{\bm{x}_{i}})=I_{\bm{x}_{i}}+s$ (with $s\in\mathbb{N}$), and then taking out $E_{I_{y_{i}}}$ as $y_{i}$. All in-context examples in a sequence share the same mapping rule $R_{s}$. To accomplish this task, the model must retrieve the same token as the query example from the context and output its subsequent token.
@@
Figure 2: The ICL test error of Transformers trained on the retrieval task with different numbers of label tokens. ’Eval’ denotes ’evaluated on’. Note that the indices of training label tokens $I_{y_{i}}\in[50,450)$, so the labels in (a) are in-distribution while (b) and (c) are OOD.
Observations. As observed in Figure 2, when the testing labels are in-distribution (left), all three models perform well. When testing labels are OOD, the more label tokens (random mappings) seen during training, the better the ICL performance is. This demonstrates that the retrieval ability can be obtained by training on a large number of similar retrieval tasks, even though the testing label tokens are unseen. This result may also provide new insights into how real-world LLMs acquire the in-context retrieval ability: by autoregressively pretraining on a huge corpus that may contain retrieval tasks, such ability emerges.
(a) PT $s\sim\mathcal{U}(50,150)$
(b) PT $s\sim\mathcal{U}(50,250)$
(c) PT $s\sim\mathcal{U}(50,450)$
Figure 3: The matching score of all attention heads of models trained on the retrieval task. ’PT’ denotes ’pretrained on’. Each subfigure corresponds to a different pretrained model. The model of (c) exhibits more heads with high matching scores, which is also the most performant model in the retrieval task in Figure 2.
The ability to perform linear regression and then retrieval can also be gained by pretraining on a similar task. To further reproduce the emergence of the abstract label learning ability of real-world LLMs, we design a task that emulates the natural language classification with abstract labels like ’foo’ and ’bar’. The task function is defined as follows: $y_{i}=f(\bm{x}_{i})=E_{I_{\bm{x}_{i}}}$, where $I_{\bm{x}_{i}}=\text{floor}(0.4*(\bm{w}^{\top}\bm{x}_{i}))+s$, with $E$ being the predefined word embedding and $s\in\mathbb{N}+$ shared in the same sequence. Here, $\bm{x}_{i}$, $\bm{w}\sim\mathcal{N}(0,1)\in\mathbb{R}^{d}$. Each $\bm{x}_{i}$ is mapped to $y_{i}$ according to $\bm{w}$ and $s$.  [^1] In our experimental setup, given a sufficiently long context ($\approx 50$), the label of the query is highly likely to appear in the context, as the number of the possible classes is far less than the number of in-context examples. In this task, estimating $\bm{w}$ and calculating $\bm{w}^{\top}\bm{x}_{i}$ simulates predicting the original label (’positive’ and ’negative’) based on the semantics in the natural language task, while retrieving the abstract labels from in-context examples that share the same $\text{floor}(0.4*(\bm{w}^{\top}\bm{x}_{i}))$ as the query from the context resembles identifying the abstract labels (’foo’ and ’bar’).
Again, we train three models on different ranges of mappings: $s\sim\mathcal{U}(100,200)$, $s\sim\mathcal{U}(100,1000)$, and $s\sim\mathcal{U}(100,2000)$, and evaluate on $s\sim\mathcal{U}(100,200)$, $s\sim\mathcal{U}(500,600)$, and $s\sim\mathcal{U}(3000,3100)$. The test error are plotted in Figure 4.
@@
Figure 4: The ICL test error of Transformers trained and tested on the linear regression + retrieval task with different numbers of label tokens. ’Eval’ denotes ’evaluated on’. Only the model trained on the largest number of tasks exhibits generalization to unseen label tokens.
Observations. In Figure 4, the generalization ability to unseen labels also improves as the number of labels encountered during training increases. Notably, only the model trained with $s\sim\mathcal{U}(100,2000)$ performs well on the unseen labels. This suggests that as long as the LLM has been exposed to sufficiently many similar tasks with during training, it can effectively address arbitrary OOD labels retrievable from context through ICL. Therefore, the ability of ICL to perform abstract label classification may not serve as an evidence of learning new tasks.
3.2 Abstract Label Classification Can Only Be Achieved on In-Distribution Tasks
A retrieval task with OOD testing functions & observations. One might question whether, once a task can be solved through retrieval, ICL can generalize beyond the training function class. To investigate this, we replace the testing functions of linear regression with quadratic regression while preserving linear regression as the pretraining task. The results in Figure 5 show that the generalization doesn’t improves with training on more in-distribution functions. Combining observations from Figure 4, we conclude that only when the testing task is in distribution can ICL solve classification with unseen labels.

### v3

@@ 1 Introduction
Recent studies have demonstrated that ICL can tackle seemingly new tasks. For instance, Garg et al. (2022); Raventós et al. (2023); Zhang et al. (2023a); Akyürek et al. (2023) found that ICL can learn new linear regression weights after pretraining on a large set of weight vectors. Moreover, Pan (2023); Kossen et al. (2024); Vacareanu et al. (2024) revealed that real-world LLMs like Llama-2 (Touvron et al., 2023) and GPT-4 (Achiam et al., 2023) are capable of solving artificially constructed tasks likely unseen in their pretraining data, such as a classification task with abstract labels.
@@
In this work, we aim to consolidate previous findings by addressing these questions. First, we empirically show that when pretrained on a specific function class, the OOD performance of ICL approaches that of a model from the same function class optimized via gradient descent. This suggests that ICL tends to implement functions encountered during pretraining, which could explain its failure on OOD tasks that significantly deviate from the training distribution. Furthermore, we reproduce the widely observed phenomenon that ICL can perform classification with abstract labels. We find that solving such tasks requires retrieving similar labels from the context, a capability that can be acquired through pretraining on analogous tasks. This implies that success in such tasks of ICL may not indicate an inherent ability to learn new tasks. Finally, we explore scenarios in which the model is pretrained on multiple tasks, empirically uncovering the algorithm selection mechanism for OOD tasks. We also provide a comprehensive theoretical framework for understanding the algorithm-selection mechanism for ICL. Our contributions are summarized as follows:
@@
    We empirically show that ICL tends to implement the pretraining function based on the downstream task context, highlighting its limitation in solving OOD tasks (Section 3.1).
@@
    We further investigate ICL’s ability to perform classification with unseen abstract labels. Although this appears to be evidence of ICL learning OOD tasks, we find that such tasks can be solved by retrieving similar examples from the context. This retrieval ability can arise from training on tasks with more diverse abstract labels (Section 4.1) and only emerges when the testing function is in distribution (Section 4.2). Additionally, we find that pretrained Llama-3-8B (Dubey et al., 2024) and Llama-2-7B fails to learn OOD functions through ICL in a synthetic word classification task (Section 4.3), further confirming ICL’s limitations in OOD scenarios.
@@
2 Existing Theoretical Predictions About ICL’s OOD Behavior
Previous literature have provided some theoretical insights into the behavior of ICL. Here we briefly review some of the representative findings. 1) ICL makes Bayesian predictions. (Xie et al., 2022; Wies et al., 2024; Zhang et al., 2023b) theoretically demonstrated that ICL behaves like a Bayesian-optimal predictor, i.e., it will infer a task concept based on the given test context, and then predict using the inferred task and the input prompt. However, these Bayesian frameworks don’t depict the concrete process of how the task concept is inferred, especially for OOD scenarios. 2) ICL implements gradient descents. Von Oswald et al. (2023); Zheng et al. (2024) construct specific Transformer weights on which the ICL prediction is equivalent to a linear regression predictor optimized by gradient descent. 3) ICL implements algorithm selection. Bai et al. (2023); Wang et al. (2024) demonstrate the existence of Transformers that can realize algorithm selection between linear classification and regression by constructing specific Transformer weights. 4) ICL performs retrieval. Li et al. (2024) proves that a trained nonlinear Transformer will concentrate its attention of the query on the in-context examples possessing similar features to that of the query.
These theoretical findings may appear disparate, as they describe different aspects of ICL under varying assumptions and settings. Furthermore, most of them remain largely unexplored empirically, particularly on large-scale nonlinear Transformers. In the following sections, we aim to provide a unifying perspective on ICL by conducting experiments with deep nonlinear Transformers on real-world OOD tasks.
3 Exploring the ICL Performance on OOD Tasks
3.1 GPT-2 Implements Functions Classes Seen During ICL Pretraining
@@
Evaluating GPT-2 on unseen mathematical function classes. To investigate the ICL performance on new tasks that are unseen during training, following Garg et al. (2022), we train a GPT-2 (Radford et al., 2019) from scratch on some simple functions and evaluate it on functions different from the training ones. Denote the Transformer model parameterized by $\theta$ as $M_{\theta}$. The pretraining objective is:
$\min_{\theta}\frac{1}{T}\sum_{i=1}^{T}\mathbb{E}_{f\sim\mathcal{F}}[\|M_{\theta}(\mathcal{S}_{i}\oplus\bm{x}_{i+1})-f(\bm{x}_{i+1})\|^{2}_{2}],$ (1)
@@
where $\mathcal{S}_{i}=[\bm{x}_{1}\oplus y_{1}\oplus\bm{x}_{2}\oplus y_{2}\oplus...\oplus\bm{x}_{i}\oplus y_{i}]\in\mathbb{R}^{d\times 2i}$ is the context of length $i$, $\oplus$ denotes concatenation. $\bm{x}_{i}\in\mathbb{R}^{d}$ are sampled from a standard Gaussian distribution $\mathcal{N}(0,1)$ with dimension $d=20$. Let $y_{i}=f(\bm{x}_{i})$ represent the labels, with $\mathcal{F}$ denoting the hypothesis class to which $f$ belongs. We train three separate GPT-2 models on three different function classes $\mathcal{F}$: linear regression (LR), quadratic regression (QR, element-wise square followed by linear regression), and a 2-layer ReLU network (ReLU NN, detailed descriptions are in Appendix C.1). We then evaluate their ICL performance on these three tasks. Note that even when the testing and training functions are i.i.d. sampled from the same task, the specific instances of the testing functions remain unseen during training. For comparison, we also train models within the corresponding $\mathcal{F}$ with gradient descent (GD) using the testing in-context examples (details in Appendix C.1).
@@
Observations. We plot the test error on the three tasks in Figure 1 and observe that: 1) (an existing finding in Garg et al. (2022)) when evaluated on the same task $\mathcal{F}$ as pretraining, ICL can reach near-zero test error. 2) (our novel finding) when evaluated on a new task, ICL performs similarly to the corresponding model of the pretraining function class optimized by GD given enough in-context examples. This indicates that the ICL prediction implements the training function classes. 3) (our novel finding) The models trained on linear and quadratic regression exhibit a double descent error curve (Nakkiran, 2019), characterized by a high error when given exact $d$ examples and evaluated on a new task. This further demonstrates that ICL implements the ID predictions, as the double descent curve is a distinctive phenomenon unique to linear regression models.
(a) Evaluated on LR
(b) Evaluated on QR
(c) Evaluated ReLU NN
@@
Figure 1: The ICL test error of Transformers trained on different function classes (solid lines) and the performance of the models from the corresponding pretraining functions classes trained by gradient descent (GD) using the in-context examples (dashed lines). Y-axis: test square error. X-axis: context length. In all evaluation tasks, we observe that as the test context length increases, the ICL performance of the Transformer pretrained on a particular function class closely approaches that of the model from this function class trained by GD.
3.2 Real-world LLMs Tend to Make In-distribution Predictions during ICL
In this section, we will demonstrate how the tendency of ICL to perform ID predictions manifests in real-world LLMs. To this end, we designed a task involving predicting labels with letters reversed. In some basic tasks like outputting antonyms or translating from English to French, all the letters of the original labels are reversed (e.g., “positive”$\rightarrow$“evitisop”). We found that in this task, a pretrained Llama-3-8B (Dubey et al., 2024) tend to output the reversed result of the query word rather than first predicting the correct label and then reversing it. Although both reversal tasks are uncommon, directly outputting the reversed version of a word is relatively more common than first reasoning and then outputting the reversed prediction. Therefore, this result reflects to some extent that LLMs, when performing ICL, are more inclined to implement ID tasks. See Appendix C.2 for more details.
(a) Antonym
(b) Country-capital
(c) English-French
(d) English-German
Figure 2: The top-1 accuracy of predicting the reversed query word (blue) and predicting the reversed target label word (orange). The accuracy of predicting the reversed query word is higher than outputting the reversed target, indicating ICL makes ID predictions.
Summary of the Empirical Results & Connections with the Existing Theories (I) Given an OOD context, ICL finds a near-optimal solution within its pretraining task space. Particularly, when learning OOD mathematical functions, ICL behaves as a predictor of its pretraining function class optimized by GD using the in-context examples. This validates and extends existing results by Zhang et al. (2023a) which theoretically shows that linear attention models trained on linear regression data still implement linear regression given arbitrary downstream context (see Appendix D).
Inspired by Raventós et al. (2023), we also explore whether increasing the diversity of training tasks, while keeping them in-distribution, can activate the OOD generalization ability. The results in Appendix B.4 also suggest a negative conclusion.
4 Learning Abstract Labels May Not Be A Real OOD Capability
@@ 4.1 Classification Tasks with Unseen Abstract Labels
Recent works (Pan, 2023; Kossen et al., 2024) have shown that LLMs can perform classification tasks in which the labels are “abstract symbols” with no semantic meaning. For instance, in the SST-2 binary classification task, the labels “positive” and “negative” are substituted with abstract terms like “foo” and “bar”, respectively. These tasks are likely not seen during pretraining. Pan (2023) refer to this ability of ICL to perform such classification as “task learning” (TL). In this section, we explore whether the TL ability really reflects a new-task-learning capability of ICL or if it merely stems from the model having learned similar tasks during pretraining.
The retrieval ability can be gained by pretraining on a retrieval task with diverse input-label mappings. The classification of abstract labels can be approached by retrieving an example with semantics similar to the query and then outputting the label of that example. Therefore, the retrieval ability is a crucial prerequisite for performing abstract-label classification. We design a retrieval task to investigate whether ICL’s retrieval capability can emerge from training on similar tasks. Specifically, we generate a predefined word embedding $E\in\mathbb{R}^{N\times d}$ and randomly sample $\bm{x}_{i}\in\mathbb{R}^{d}$ from the first 5 rows of $E$. Each vector $\bm{x}_{i}$ corresponds to the $I_{\bm{x}_{i}}$-th row of $E$, i.e., $\bm{x}_{i}=E_{I_{\bm{x}_{i}}}$. To generate the labels $y_{i}$, we follow these steps: First, map the index $I_{\bm{x}_{i}}$ to a new one $I_{y_{i}}\in[N]$ using the mapping rule $I_{y_{i}}=I_{\bm{x}_{i}}+s$, where $s\in\mathbb{N}$ is randomly sampled. Second, we set $y_{i}=E_{I_{y_{i}}}$. All in-context examples in a sequence share the same mapping rule defined by $s$. To succeed in this task, the model must retrieve the same token as the query example from the context and output its subsequent token. All models are trained with 200,000$\times$64 sequences, where 200,000 is the number of training steps and 64 is the batch size.
@@
Figure 3: The ICL test error of Transformers trained on the retrieval task with different numbers of label tokens. “Eval” denotes “evaluated on”. Note that the indices of training label tokens $I_{y_{i}}\in[50,455)$, so the labels in (a) are ID while (b) and (c) are OOD.
Observations. In Figure 3, all three models perform well when labels are in distribution (a). When the labels are OOD, the ICL performance improves with the number of label vectors (random mappings) encountered during training. This demonstrates that retrieval ability can emerge from training on diverse retrieval tasks. These findings may also offer new insights into how real-world LLMs develop in-context retrieval capabilities: when autoregressive pretraining includes numerous instances requiring the model to retrieve tokens from previous contexts, such abilities can emerge. We further validate this finding by observing the transition of the attention maps in Appendix B.1.
The ability to perform linear regression and then retrieval can also be gained by pretraining on a similar task. To further reproduce the emergence of the abstract label learning ability of real-world LLMs, we design a task that emulates the natural language classification with abstract labels like “foo” and “bar”. The task function is defined as follows: $y_{i}=f(\bm{x}_{i})=E_{I_{\bm{x}_{i}}}$, where $I_{\bm{x}_{i}}=\text{floor}(0.4*(\bm{w}^{\top}\bm{x}_{i}))+s$, with $E$ being the predefined word embedding and $s\in\mathbb{N}+$ shared in the same sequence. Here, $\bm{x}_{i}$, $\bm{w}\sim\mathcal{N}(0,1)\in\mathbb{R}^{d}$. Each $\bm{x}_{i}$ is mapped to $y_{i}$ according to $\bm{w}$ and $s$.  [^1] In our experimental setup, given a sufficiently long context ($\approx 50$), the label of the query is highly likely to appear in the context, as the number of the possible classes is far less than the number of in-context examples. In this task, estimating $\bm{w}$ and calculating $\bm{w}^{\top}\bm{x}_{i}$ simulates predicting the original label (“positive” and “negative”) based on the semantics in the natural language task, while retrieving the abstract labels from in-context examples that share the same $\text{floor}(0.4*(\bm{w}^{\top}\bm{x}_{i}))$ as the query from the context resembles identifying the abstract labels (“foo” and “bar”).
Again, we train three models on different ranges of mappings: $s\sim\mathcal{U}(100,200)$, $s\sim\mathcal{U}(100,1000)$, and $s\sim\mathcal{U}(100,2000)$, and evaluate on $s\sim\mathcal{U}(100,200)$, $s\sim\mathcal{U}(500,600)$, and $s\sim\mathcal{U}(3000,3100)$. The test error is plotted in Figure 4.
@@
Figure 4: The ICL test error of Transformers trained and tested on the linear regression + retrieval task with different numbers of label tokens. “Eval” denotes “evaluated on”. Only the model trained on the largest number of tasks exhibits generalization to unseen label tokens.
Observations. In Figure 4, the generalization ability to unseen labels also improves as the number of labels encountered during training increases. Notably, only the model trained with $s\sim\mathcal{U}(100,2000)$ performs well on the unseen labels. This suggests that as long as the LLM has been exposed to sufficiently many similar tasks during training, it can effectively classify arbitrary OOD labels retrievable from context through ICL.
4.2 Abstract Label Classification Can Only Be Achieved on ID Tasks
A retrieval task with OOD testing functions & observations. Given the above observations, one might question whether, once the target labels appear in the context, ICL can generalize beyond the training function class by retrieving the target label from the context. To investigate this, we conduct the same predict-then-retrieval task as in Figure 4 but replace the testing functions with quadratic regression while preserving linear regression as the pretraining task. The results in Figure 5 show that the generalization doesn’t improve with training on more ID functions. Combining observations from Figure 4, we conclude that ICL can only solve classification with unseen labels over ID test function classes. Once the underlying task function is OOD, ICL fails even if the target label appears in the context. This finding highlights a limitation in improving an LLM’s performance through in-context examples. While providing examples with shared labels may seem helpful, this approach may fail if the underlying prediction rule is too OOD for the LLM to learn.